# Orbital optimized vs time-dependent density functional calculations of intramolecular charge transfer excited states

### v3

@@ Abstract
The performance of time-independent, orbital optimized calculations of excited states is assessed with respect to charge transfer excitations in organic molecules in comparison to the linear-response time-dependent density functional theory (TD-DFT) approach. A direct optimization method to converge on saddle points of the electronic energy surface is used to carry out calculations with the local density approximation (LDA) and the generalized gradient approximation (GGA) functionals PBE and BLYP for a set of 27 excitations in 15 molecules. The time-independent approach is fully variational and provides a relaxed excited state electron density from which the extent of charge transfer is quantified. The TD-DFT calculations are generally found to provide larger charge transfer distances compared to the orbital optimized calculations, even when including first-order orbital relaxation effects with the Z-vector method. While the error on the excitation energy relative to theoretical best estimates is found to increase with the extent of charge transfer up to ca. eV for TD-DFT, no correlation is observed for the orbital optimized approach. The orbital optimized calculations with the LDA and the GGA functionals provide a mean absolute error of 0.7 eV, outperforming TD-DFT with both local and global hybrid functionals for excitations with long-range charge transfer character. Orbital optimized calculations with the global hybrid functional B3LYP and the range-separated hybrid functional CAM-B3LYP on a selection of states with short- and long-range charge transfer indicate that inclusion of exact exchange has a small effect on the charge transfer distance, while it significantly improves the excitation energy, with the best performing functional CAM-B3LYP providing an absolute error typically around 0.15 eV.
@@ 2.1.1 Freeze-and-release direct optimization
Typically, the initial guess for a calculation of a single-electron excitation is constructed by performing a 90 rotation between an occupied orbital and an unoccupied orbital of the ground state corresponding to swapping the occupation numbers of the two orbitals. For example, for a LUMOHOMO excitation, a 90 rotation between the HOMO and the LUMO is performed, meaning that in the initial guess a hole is left in the ground state HOMO (occupation number of 0) and an electron is placed in the ground state LUMO (occupation number of 1). In the DO-MOM method, MOM is employed to restore the non-aufbau occupation of the orbitals in case of variational collapse during the electronic optimization. However, this approach has been shown not to be able to prevent variational collapse to unphysical, charge-delocalized solutions for some intramolecular charge transfer excitations49.
@@ 2.2 Computational settings
where is the difference between the real space excited and ground state electron densities, is the transferred charge defined as the integral over all space of the positive part of , and corresponds to the difference between the electric dipole moments of the ground and excited states. A charge transfer distance proportional to the difference between dipole moments is a better measure of the extent of charge transfer than the dipole moment of the excited state, since the movement of charge can be from a region with excess negative charge to a region with excess positive charge, thereby lowering the dipole moment of the molecule. Since represents the distance between the centroids of the positive and negative parts of the electron density difference, Le Bahers’ metric defines the distance of a unidirectional displacement of charge. The excitations considered in the present work have a prevalent unidirectional character, in which case the charge transfer distance computed according to eq 6 has been found to be in good agreement with more advanced metrics, which take the density shift in all directions into account15. For the time-independent calculations, the difference density, , is calculated as the difference between the fully relaxed electron density of the spin-mixed excited state solution and the ground state density. For linear-response TD-DFT, the one-particle difference density matrix from which is obtained consists of two parts95, 96. The first part, often called unrelaxed difference density matrix, is obtained from the eigenvectors of the Casida equations and includes only occupied-occupied and virtual-virtual orbital contributions. The second part contains occupied-unoccupied orbital contributions to first order and is obtained by solving the so-called Z-vector equations97, 98. In the present work, the electron density obtained from the full TD-DFT one-particle difference density matrix is referred to as relaxed difference density, as is commonly done9, 95, 96.
@@ 2.3 Characterizing and selecting multiple orbital optimized solutions
Due to the non-linearity of the orbital optimization, orbital optimized calculations can be affected by the presence of stationary points on the electronic energy surface without a clear association with a particular electronic state76, 59, 75. In the context of intramolecular charge transfer excitations, it has been shown49 that orbital optimized calculations can converge on solutions with significantly different degree of charge transfer depending on the method used. This is exemplified by the A LUMO+1HOMO charge transfer excitation in the twisted PP molecule49. A calculation of the spin-mixed solution for the A excited state of twisted PP using the PBE functional is initialized by promoting an electron from the ground state HOMO localized on the pyrrole group to the ground state LUMO+1 localized on the phenyl group. A calculation with DO and MOM converges to a solution corresponding to a 2nd-order saddle point with an excitation energy of 4.61 eV. The freeze-and-release DO approach, instead, converges to a higher-energy solution corresponding to a 10th-order saddle point with an excitation energy of 5.56 eV (5.61 eV after spin purification using eq 5), much closer to the theoretical best estimate of 5.65 eV 9. Figure 1 shows the electron density difference between ground and excited states, for the two solutions of the A excitation in the twisted PP molecule.
The solution obtained with freeze-and-release DO has larger dipole moment and charge transfer distance (9.36 D and 2.42 Å) compared to the solution obtained with DO-MOM (3.33 D and 2.06 Å). The dipole moment and charge transfer distance of the more charge-localized solution are in good agreement with the dipole moment and charge transfer distance obtained from the relaxed density of a TD-DFT calculation with the range-separated CAM-B3LYP functional (9.56 D and 2.35 Å), a functional that is known to provide an accurate description of intramolecular charge transfer excitations within TD-DFT16, 9, 19. Clearly, the orbital optimized solution that best describes the target charge transfer excited state is the one obtained in the freeze-and-release DO calculations, while DO-MOM collapses to a solution where the transferred charge is too delocalized. This delocalization is caused by mixing of the hole originally localized on the pyrrole group with a lower-energy occupied orbital initially localized on the phenyl group49.
@@ 3 Results
| Aminobenzonitrile (ABN) | A | 5.09 | 4.65 | 4.61 | 4.65 | 3.97 | 3.98 | 4.01 | 1.06 | 0.03 |
| Aniline | A | 5.48 | 5.15 | 5.07 | 5.13 | 4.66 | 4.67 | 4.72 | 0.82 | 0.02 |
| Azulene | A | 3.84 | 3.48 | 3.43 | 3.47 | 3.32 | 3.35 | 3.39 | 0.94 | 0.03 |
| B | 4.49 | 4.45 | 4.41 | 4.45 | 4.17 | 4.09 | 4.12 | 0.77 | 0.03 | |
| Benzonitrile | A | 7.05 | 5.75 | 5.75 | 5.76 | 6.87 | 6.06 | 6.66 | 1.04 | 0.49 |
| Benzothiadiazole (BTD) | B | 4.28 | 3.58 | 3.52 | 3.57 | 3.41 | 3.42 | 3.46 | 1.19 | 0.04 |
| Dimethylaminobenzonitrile (DMABN) | A | 4.86 | 4.31 | 4.32 | 4.34 | 3.77 | 3.79 | 3.80 | 1.56 | 0.05 |
| Twisted DMABN | A | 4.12 | 2.17 | 2.36 | 2.30 | 3.59 | 3.62 | 3.55 | 2.04 | 0.12 |
| B | 4.75 | 2.76 | 2.95 | 2.90 | 4.25 | 4.30 | 4.23 | 1.75 | 0.12 | |
| Dimethylaniline (DMAn) | B | 4.40 | 3.94 | 4.01 | 4.01 | 3.98 | 3.97 | 3.99 | 1.08 | 0.06 |
| A | 5.40 | 4.89 | 4.93 | 4.97 | 4.46 | 4.49 | 4.52 | 1.33 | 0.04 | |
@@
| Nitroaniline | A | 4.39 | 3.50 | 3.48 | 3.53 | 3.46 | 3.45 | 3.49 | 2.05 | 0.16 |
| Nitrobenzene | A | 5.39 | 4.36 | 4.31 | 4.38 | 4.45 | 4.45 | 4.53 | 1.46 | 0.11 |
| Nitrodimethylaniline (NDMA) | A | 4.13 | 3.19 | 3.20 | 3.23 | 3.21 | 3.21 | 3.24 | 2.34 | 0.18 |
| Nitropyridine -Oxide (NPNO) | A | 4.10 | 3.47 | 3.43 | 3.48 | 3.05 | 3.03 | 3.08 | 1.72 | 0.13 |
| -Phenylpyrrole (PP) | B | 5.32 | 4.01 | 4.01 | 4.03 | 4.41 | 4.45 | 4.51 | 1.59 | 0.20 |
| A | 5.86 | 3.93 | 3.94 | 3.96 | 5.34 | 5.29 | 5.32 | 2.02 | 0.35 | |
| Twisted PP | B | 5.58 | 3.56 | 3.58 | 3.59 | 5.41 | 5.29 | 5.29 | 2.36 | 0.15 |
| A | 5.65 | 3.60 | 3.62 | 3.63 | 5.71 | 5.59 | 5.61 | 2.41 | 0.27 | |
| A | 5.95 | 4.16 | 4.17 | 4.19 | 5.50 | 5.43 | 5.41 | 2.15 | 0.21 | |
| B | 6.17 | 4.19 | 4.21 | 4.22 | 5.57 | 5.59 | 5.59 | 2.16 | 0.27 | |
| Phthalazine | A | 3.91 | 2.50 | 2.71 | 2.66 | 3.17 | 3.22 | 3.17 | 1.26 | 0.11 |
| B | 4.31 | 3.10 | 3.29 | 3.25 | 3.66 | 3.68 | 3.64 | 1.26 | 0.07 | |
| Quinoxaline | B | 4.63 | 3.73 | 3.71 | 3.75 | 3.78 | 3.81 | 3.85 | 1.25 | 0.05 |
| A | 5.65 | 5.43 | 5.42 | 5.47 | 4.75 | 4.66 | 4.69 | 0.62 | 0.04 | |
| B | 6.22 | 4.67 | 4.82 | 4.78 | 5.20 | 5.25 | 5.19 | 1.24 | 0.13 | |
@@
The electronic distance from the initial guess is typically between 0.02-0.1 for states with and between 0.1-0.3 for states with . The distance from the initial guess systematically increases with the charge transfer distance because the orbital relaxation effects are more pronounced. However, there are two cases where the distance from the initial guess is larger than expected based on this trend, the A state of benzonitrile and the A state of planar PP. For LDA and PBE orbital optimized calculations with an initial guess targeting the A excited state of benzonitrile, two spin-mixed solutions are found. In the case of PBE, the first solution, obtained with DO-MOM, has and an excitation energy after spin purification of 6.02 eV. The charge transfer distance and dipole moment of this solution are 0.58 Å and 2.81 D, respectively. The second solution, obtained with freeze-and-release DO, is farther from the initial guess (), has an excitation energy of 6.66 eV, and charge transfer distance and dipole moment of 1.04 Å and 0.58 D, respectively (note that the charge transfer connected to this excitation lowers the dipole moment of the molecule). The theoretical best estimate of the excitation energy is 7.05 eV9, while previous TD-DFT calculations with the range-separated hybrid functional CAM-B3LYP and CASSCF calculations101 report a charge transfer distance and dipole moment of 1.17 Å and 0.64 D, respectively. Therefore, in this case it appears that the solution with largest distance from the initial guess is the one best describing the target charge transfer excitation. In the case of the A excited state of planar PP ( when using the PBE functional), the spin-mixed solution obtained with any of the three functionals exhibits greater delocalization of the transferred charge than the solution obtained for the A state at the twisted geometry (see section 2.3), despite previous calculations indicating a small dependence of the strength of charge transfer on the geometry9. Any attempts using freeze-and-release DO and a DO approach based on generalized mode following (DO-GMF)49 could not find a spin-mixed excited state solution with enhanced charge localization closer to the initial guess for this state.
@@
| Azulene | B | 4.49 | 4.63 | 0.92 | 4.77 | 1.02 | 4.38 | 0.80 | 4.71 | 0.85 |
| Quinoxaline | B | 6.22 | 5.69 | 1.21 | 6.47 | 1.24 | 5.82 | 1.22 | 6.34 | 1.22 |
| Twisted DMABN | B | 4.75 | 3.84 | 1.79 | 4.69 | 1.74 | 4.49 | 1.70 | 4.60 | 1.70 |
| NDMA | A | 4.13 | 3.66 | 2.32 | 4.13 | 2.26 | 3.57 | 2.28 | 4.01 | 2.28 |
| Twisted PP | A | 5.95 | 4.96 | 2.49 | 5.89 | 2.35 | 5.66 | 2.29 | 5.80 | 2.37 |
@@ 3.0.1 Charge transfer distance
In most cases, the functional has a small effect on the charge transfer distance, especially for the orbital optimized calculations. Apart from the state with shortest charge transfer distance (B state of azulene), the orbital optimized calculations using the range-separated hybrid functional CAM-B3LYP give charge transfer distances very close to those obtained from the TD-DFT calculations using CAM-B3LYP.

### v4

# Orbital-optimized versus time-dependent density functional calculations of intramolecular charge transfer excited states

@@ Abstract
The performance of time-independent, orbital optimized calculations of excited states is assessed with respect to charge transfer excitations in organic molecules in comparison to the linear-response time-dependent density functional theory (TD-DFT) approach. A direct optimization method to converge on saddle points of the electronic energy surface is used to carry out calculations with the local density approximation (LDA) and the generalized gradient approximation (GGA) functionals PBE and BLYP for a set of 27 excitations in 15 molecules. The time-independent approach is fully variational and provides a relaxed excited state electron density from which the extent of charge transfer is quantified. The TD-DFT calculations are generally found to provide larger charge transfer distances compared to the orbital optimized calculations, even when including orbital relaxation effects with the Z-vector method. While the error on the excitation energy relative to theoretical best estimates is found to increase with the extent of charge transfer up to ca. eV for TD-DFT, no correlation is observed for the orbital optimized approach. The orbital optimized calculations with the LDA and the GGA functionals provide a mean absolute error of 0.7 eV, outperforming TD-DFT with both local and global hybrid functionals for excitations with long-range charge transfer character. Orbital optimized calculations with the global hybrid functional B3LYP and the range-separated hybrid functional CAM-B3LYP on a selection of states with short- and long-range charge transfer indicate that inclusion of exact exchange has a small effect on the charge transfer distance, while it significantly improves the excitation energy, with the best performing functional CAM-B3LYP providing an absolute error typically around 0.15 eV.
@@ 2.1.1 Freeze-and-release direct optimization
Typically, the initial guess for a calculation of a single-electron excitation is constructed by performing a 90∘ rotation between an occupied orbital and an unoccupied orbital of the ground state corresponding to swapping the occupation numbers of the two orbitals. For example, for a LUMOHOMO excitation, a 90∘ rotation between the HOMO and the LUMO is performed, meaning that in the initial guess a hole is left in the ground state HOMO (occupation number of 0) and an electron is placed in the ground state LUMO (occupation number of 1). In the DO-MOM method, MOM is employed to restore the non-aufbau occupation of the orbitals in case of variational collapse during the electronic optimization. However, this approach has been shown not to be able to prevent variational collapse to unphysical, charge-delocalized solutions for some intramolecular charge transfer excitations49.
@@ 2.2 Computational settings
where is the difference between the real space excited and ground state electron densities, is the transferred charge defined as the integral over all space of the positive part of , and corresponds to the difference between the electric dipole moments of the ground and excited states. A charge transfer distance proportional to the difference between dipole moments is a better measure of the extent of charge transfer than the dipole moment of the excited state, since the movement of charge can be from a region with excess negative charge to a region with excess positive charge, thereby lowering the dipole moment of the molecule. Since represents the distance between the centroids of the positive and negative parts of the electron density difference, Le Bahers’ metric defines the distance of a unidirectional displacement of charge. The excitations considered in the present work have a prevalent unidirectional character, in which case the charge transfer distance computed according to eq 6 has been found to be in good agreement with more advanced metrics, which take the density shift in all directions into account15. For the time-independent calculations, the difference density, , is calculated as the difference between the fully relaxed electron density of the spin-mixed excited state solution and the ground state density. For linear-response TD-DFT, the one-particle difference density matrix from which is obtained consists of two parts95, 96. The first part, often called unrelaxed difference density matrix, is obtained from the eigenvectors of the Casida equations and includes only occupied-occupied and virtual-virtual orbital contributions. The second part contains occupied-unoccupied orbital contributions and is obtained by solving the so-called Z-vector equations97, 98. In the present work, the electron density obtained from the full TD-DFT one-particle difference density matrix is referred to as relaxed difference density, as is commonly done9, 95, 96.
@@ 2.3 Characterizing and selecting multiple orbital optimized solutions
Due to the non-linearity of the orbital optimization, orbital optimized calculations can be affected by the presence of stationary points on the electronic energy surface without a clear association with a particular electronic state76, 59, 75. In the context of intramolecular charge transfer excitations, it has been shown49 that orbital optimized calculations can converge on solutions with significantly different degree of charge transfer depending on the method used. This is exemplified by the A1 LUMO+1HOMO charge transfer excitation in the twisted PP molecule49. A calculation of the spin-mixed solution for the A1 excited state of twisted PP using the PBE functional is initialized by promoting an electron from the ground state HOMO localized on the pyrrole group to the ground state LUMO+1 localized on the phenyl group. A calculation with DO and MOM converges to a solution corresponding to a 2nd-order saddle point with an excitation energy of 4.61 eV. The freeze-and-release DO approach, instead, converges to a higher-energy solution corresponding to a 10th-order saddle point with an excitation energy of 5.56 eV (5.61 eV after spin purification using eq 5), much closer to the theoretical best estimate of 5.65 eV 9. Figure 1 shows the electron density difference between ground and excited states, for the two solutions of the A1 excitation in the twisted PP molecule.
The solution obtained with freeze-and-release DO has larger dipole moment and charge transfer distance (9.36 D and 2.42 Å) compared to the solution obtained with DO-MOM (3.33 D and 2.06 Å). The dipole moment and charge transfer distance of the more charge-localized solution are in good agreement with the dipole moment and charge transfer distance obtained from the relaxed density of a TD-DFT calculation with the range-separated CAM-B3LYP functional (10.16 D and 2.41 Å), a functional that is known to provide an accurate description of intramolecular charge transfer excitations within TD-DFT16, 9, 19. Clearly, the orbital optimized solution that best describes the target charge transfer excited state is the one obtained in the freeze-and-release DO calculations, while DO-MOM collapses to a solution where the transferred charge is too delocalized. This delocalization is caused by mixing of the hole originally localized on the pyrrole group with a lower-energy occupied orbital initially localized on the phenyl group49.
@@ 3 Results
| Aminobenzonitrile (ABN) | A1 | 5.09 | 4.65 | 4.61 | 4.65 | 3.97 | 3.98 | 4.01 | 1.06 | 0.03 |
| Aniline | A1 | 5.48 | 5.15 | 5.07 | 5.13 | 4.66 | 4.67 | 4.72 | 0.82 | 0.02 |
| Azulene | A1 | 3.84 | 3.48 | 3.43 | 3.47 | 3.32 | 3.35 | 3.39 | 0.94 | 0.03 |
| B2 | 4.49 | 4.45 | 4.41 | 4.45 | 4.17 | 4.09 | 4.12 | 0.77 | 0.03 | |
| Benzonitrile | A2 | 7.05 | 5.75 | 5.75 | 5.76 | 6.87 | 6.06 | 6.66 | 1.04 | 0.49 |
| Benzothiadiazole (BTD) | B2 | 4.28 | 3.58 | 3.52 | 3.57 | 3.41 | 3.42 | 3.46 | 1.19 | 0.04 |
| Dimethylaminobenzonitrile (DMABN) | A1 | 4.86 | 4.31 | 4.32 | 4.34 | 3.77 | 3.79 | 3.80 | 1.56 | 0.05 |
| Twisted DMABN | A2 | 4.12 | 2.17 | 2.36 | 2.30 | 3.59 | 3.62 | 3.55 | 2.04 | 0.12 |
| B1 | 4.75 | 2.76 | 2.95 | 2.90 | 4.25 | 4.30 | 4.23 | 1.75 | 0.12 | |
| Dimethylaniline (DMAn) | B2 | 4.40 | 3.94 | 4.01 | 4.01 | 3.98 | 3.97 | 3.99 | 1.08 | 0.06 |
| A1 | 5.40 | 4.89 | 4.93 | 4.97 | 4.46 | 4.49 | 4.52 | 1.33 | 0.04 | |
@@
| Nitroaniline | A1 | 4.39 | 3.50 | 3.48 | 3.53 | 3.46 | 3.45 | 3.49 | 2.05 | 0.16 |
| Nitrobenzene | A1 | 5.39 | 4.36 | 4.31 | 4.38 | 4.45 | 4.45 | 4.53 | 1.46 | 0.11 |
| Nitrodimethylaniline (NDMA) | A1 | 4.13 | 3.19 | 3.20 | 3.23 | 3.21 | 3.21 | 3.24 | 2.34 | 0.18 |
| Nitropyridine -Oxide (NPNO) | A1 | 4.10 | 3.47 | 3.43 | 3.48 | 3.05 | 3.03 | 3.08 | 1.72 | 0.13 |
| -Phenylpyrrole (PP) | B2 | 5.32 | 4.01 | 4.01 | 4.03 | 4.41 | 4.45 | 4.51 | 1.59 | 0.20 |
| A1 | 5.86 | 3.93 | 3.94 | 3.96 | 5.34 | 5.29 | 5.32 | 2.02 | 0.35 | |
| Twisted PP | B2 | 5.58 | 3.56 | 3.58 | 3.59 | 5.41 | 5.29 | 5.29 | 2.36 | 0.15 |
| A1 | 5.65 | 3.60 | 3.62 | 3.63 | 5.71 | 5.59 | 5.61 | 2.41 | 0.27 | |
| A2 | 5.95 | 4.16 | 4.17 | 4.19 | 5.50 | 5.43 | 5.41 | 2.15 | 0.21 | |
| B1 | 6.17 | 4.19 | 4.21 | 4.22 | 5.57 | 5.59 | 5.59 | 2.16 | 0.27 | |
| Phthalazine | A2 | 3.91 | 2.50 | 2.71 | 2.66 | 3.17 | 3.22 | 3.17 | 1.26 | 0.11 |
| B1 | 4.31 | 3.10 | 3.29 | 3.25 | 3.66 | 3.68 | 3.64 | 1.26 | 0.07 | |
| Quinoxaline | B2 | 4.63 | 3.73 | 3.71 | 3.75 | 3.78 | 3.81 | 3.85 | 1.25 | 0.05 |
| A1 | 5.65 | 5.43 | 5.42 | 5.47 | 4.75 | 4.66 | 4.69 | 0.62 | 0.04 | |
| B1 | 6.22 | 4.67 | 4.82 | 4.78 | 5.20 | 5.25 | 5.19 | 1.24 | 0.13 | |
@@
The electronic distance from the initial guess is typically between 0.02-0.1 for states with and between 0.1-0.3 for states with . The distance from the initial guess systematically increases with the charge transfer distance because the orbital relaxation effects are more pronounced. However, there are two cases where the distance from the initial guess is larger than expected based on this trend, the A2 state of benzonitrile and the A1 state of planar PP. For LDA and PBE orbital optimized calculations with an initial guess targeting the A2 excited state of benzonitrile, two spin-mixed solutions are found. In the case of PBE, the first solution, obtained with DO-MOM, has and an excitation energy after spin purification of 6.02 eV. The charge transfer distance and dipole moment of this solution are 0.58 Å and 2.81 D, respectively. The second solution, obtained with freeze-and-release DO, is farther from the initial guess (), has an excitation energy of 6.66 eV, and charge transfer distance and dipole moment of 1.04 Å and 0.58 D, respectively (note that the charge transfer connected to this excitation lowers the dipole moment of the molecule). The theoretical best estimate of the excitation energy is 7.05 eV9, while previous TD-DFT calculations with the range-separated hybrid functional CAM-B3LYP and CASSCF calculations101 report a charge transfer distance and dipole moment of 1.17 Å and 0.64 D, respectively. Therefore, in this case it appears that the solution with largest distance from the initial guess is the one best describing the target charge transfer excitation. In the case of the A1 excited state of planar PP ( when using the PBE functional), the spin-mixed solution obtained with any of the three functionals exhibits greater delocalization of the transferred charge than the solution obtained for the A1 state at the twisted geometry (see section 2.3), despite previous calculations indicating a small dependence of the strength of charge transfer on the geometry9. Any attempts using freeze-and-release DO and a DO approach based on generalized mode following (DO-GMF)49 could not find a spin-mixed excited state solution with enhanced charge localization closer to the initial guess for this state.
@@
| Azulene | B2 | 4.49 | 4.63 | 0.92 | 4.77 | 1.02 | 4.38 | 0.80 | 4.71 | 0.85 |
| Quinoxaline | B1 | 6.22 | 5.69 | 1.21 | 6.47 | 1.24 | 5.82 | 1.22 | 6.34 | 1.22 |
| Twisted DMABN | B1 | 4.75 | 3.84 | 1.79 | 4.69 | 1.74 | 4.49 | 1.70 | 4.60 | 1.70 |
| NDMA | A1 | 4.13 | 3.66 | 2.32 | 4.13 | 2.26 | 3.57 | 2.28 | 4.01 | 2.28 |
| Twisted PP | A2 | 5.95 | 4.96 | 2.49 | 5.89 | 2.35 | 5.66 | 2.29 | 5.80 | 2.37 |
@@ 3.0.1 Charge transfer distance
In most cases, the functional has a small effect on the charge transfer distance, especially for the orbital optimized calculations. Apart from the state with shortest charge transfer distance (B2 state of azulene), the orbital optimized calculations using the range-separated hybrid functional CAM-B3LYP give charge transfer distances very close to those obtained from the TD-DFT calculations using CAM-B3LYP.